Installing a muffler at Barut Garage, Inc. involves a five work tasks. Azra has timed workers performing these tasks seven times, with the results (in minutes) shown in the following table.

No

Tasks

1

2

3

4

5

6

7

PR

1

Select correct mufflers

5

4

6

3

4

5

5

120

2

Remove old muffler

8

8

8

6

8

7

6

85

3

Install new muffler

15

13

15

14

12

15

13

85

4

Inspect the work

4

4

3

3

3

3

3

110

5

Complete paperwork

9

8

9

8

10

7

10

110


  
      
Given that Azra does not want an error of more than 0.07 and a confidence level less than 0.85 answer the following:
  
Provide the z-score:  Round to the nearest hundredth, for example 4.22
  
Does the number of observations for the following tasks warrant such confidence level?  
Enter YES or NO.
  
Task # 3 Install new muffler      :  
Task # 5 Complete paperwork   :  
  
What are the required numbers of observations for these tasks? Enter the number.
  
Task # 3 Install new muffler      :  
Task # 5 Complete paperwork   :  
  
Assume that the number of observations for the following tasks justifies the confidence level and accuracy mentioned above. Also note that Azra factors in a total of fatigue and personal allowance of 0.15 applied to job time. Provide the following results from your solution:

Task

Normal Time*

Standard Time*

# 2 Remove old muffler

min

min

# 4 Inspect the work

min

min


  
Provide the total standard time* for the whole job:   minutes.
*Round to the nearest hundredth, for example 8.45 minutes
Answer #1
Task 1 2 3 4 5 6 7 PR mean NT
1 5 4 6 3 4 5 5 120 4.571429 5.485714
2 8 8 8 6 8 7 6 85 7.285714 6.192857
3 15 13 15 14 12 15 13 85 13.85714 11.77857
4 4 4 3 3 3 3 3 110 3.285714 3.614286
5 9 8 9 8 10 7 10 110 8.714286 9.585714

for 0.85 confidence level, z score will be 0.935

Number of observations needed for a task =[ z xSD / h x Mean]2

where h is the tolerance level

for task 3 N= [ 0.935 x1.2149 / 0.07x13.857]2 = 1.37 =2

hence the process does not warrant such high confidence level,since current number of observations is 7

For task 5 N = [ 0.935 x 1.1126 / 0.07x8.7142]2 = 2.9

Hence the process does not warrant such high confidence level, since current number of observations is 7

Task 2 normal time = rating * mean time = 6.1928

Task 2 standard time = 6.1928 /1- fatigue allowance = 6.1928 /1-0.15 =7.2856

Task 4 normal time = 3.6142

Task 4 standard time = 3,6142 / 1- 0.15 =4.252

Normal time for complete task = 36.657

Standard time for complete task = 36.657 /1-0.15 = 43.1258